OpenStudy (anonymous):

Solve by completing the square. x^2 +6x-27=0 x=

OpenStudy (anonymous):

x^2+6x=27 x^2+6x+(___)^2=27+(_____)^2 So in the (_____), you take the number in front of the x (not x^2, just x), then divide it by 2. Then you would square it. In this case it would be 6/2, which equals 3.. then you square it = 9. In cases where you have another number in front of x^2 (besides 1), you need to make it 1. In this case, the number in front of x^2 is already 1, so you don't need to do anything to make it 1. x^2+6x+(9)=27+(9). (x+3)^2 = 36 x+3 = 6 or x+3= -6 Don't forget the positive and negative that is associated with radicals. x=?

OpenStudy (anonymous):

Using the reverse of the foil method x^2 +6x-27=(x+9)(x-3) so x= -9 or x=3

jhonyy9 (jhonyy9):

x_1 = -9 and x_2 = 3 are right solutions

OpenStudy (anonymous):

x^2+2(3)x+9-9-27=0 (x+3)^2-36=0 (x+3)^2-6^2=0 (x+3+6)(x+3-6)=0 x=3,-9

OpenStudy (anonymous):

x^2+6x-27=0